# Decomposing Multivariate Information Rates in Networks of Random Processes

### v2

@@ V.2 Data and statistical analysis
The four physiological time series described above were taken as a realization of the random processes F, M, H and R. These processes were investigated to decompose the rate of information shared dynamically between the target $Y=F$ and the set of source processes $X=\{H,M,R\}$, with $H=X_{1},M=X_{2},R=X_{3}$, with the aim of investigating the mechanisms underlying the short-term dynamic regulation of the cerebral blood flow in resting and stress conditions. Each series was first detrended with an AR high-pass filter with zero phase (cutoff frequency 0.015 cycles/beat) Nollo et al. (2000). Then, a VAR model in the form of (29) was fitted to the four time series; model identification was performed via the ordinary least-squares approach Lütkepohl (2005), setting the model order $p$ according to the Akaike Information Criterion (AIC) for each subject (with maximum model order equal to 14) Akaike (1974b). After VAR identification, computation of time and frequency domain interaction measures of mutual, unique, redundant and synergistic information rates was performed from the estimated model parameters and spectra of the processes. Spectral analysis was performed assuming the series as uniformly sampled with the mean HP taken as the sampling period ($f_{s}=\frac{1}{<H_{n}>}$). Specifically, the spectral MIRs between the target and groups of sources were computed as in (26); then, according to coarse-graining PIRD, the unique, redundant and synergistic information rates calculated from the spectral redundancy rate in (21) were computed through integration within the low frequency (LF, $\left[0.04-0.15\right]$ Hz) and high frequency (HF, $\left[0.15-0.4\right]$ Hz) bands of the spectrum to analyze specific rhythms with physiological meaning Malpas (2002), as well as through integration over all frequencies, to get overall time-domain measures.
@@
Figure 6: Coarse-grained PIRD applied to the physiological network of mean cerebral blood flow velocity ($F$), mean arterial pressure ($M$), heart period ($H$) and respiration ($R$) assessed in patients prone to develop postural-related syncope. The mutual (left column: $I_{F;H}$ (black dots), $I_{F;M}$ (gray dots), $I_{F;R}$ (pink dots), $I_{F;H,M,R}$ (light blue dots)), unique (middle column: $\mathcal{U}_{F;H}$ (black dots), $\mathcal{U}_{F;M}$ (gray dots), $\mathcal{U}_{F;R}$ (pink dots)), redundant (right column: $\mathcal{R}_{F;H;M;R}$, magenta dots) and synergistic (right column: $\mathcal{S}_{F;H;M;R}$, blue dots) information rates shared between the target $F$ and the sources $\{H,M,R\}$ are computed in the RS (left boxplots), ET (middle boxplots) and LT (right boxplots) conditions along a) the whole frequency axis (TIME), b) the LF and c) the HF bands of the spectrum, taking the spectral redundancy rate function as in (21). The redundancy-synergy balance $\Delta$ is shown as boxplot distributions and individual values (purple dots) in the right column of panels a) (TIME), b) (LF) and c) (HF). The statistical differences among pairs of unique information rates (middle panels) were tested via Wilcoxon signed rank test for paired data ($p<0.05$) with Bonferroni-Holm correction for multiple comparisons: significant p-values are shown above the boxplot distributions with colors indicating the distribution to be compared with (i.e., black if compared with $\mathcal{U}_{F;H}$, gray if compared with $\mathcal{U}_{F;M}$). Wilcoxon signed rank test for paired data, $p<0.05$: $\ast$, RS vs. ET; §$\S$, RS vs. LT; ##\##, ET vs. LT.
@@ V.3 Results and discussion
Overall, these results confirm the redundant nature of cardiovascular and cerebrovascular interactions previously reported for similar triplets of physiological processes Faes et al. (2017); Porta et al. (2017); Faes et al. (2022), and document the relevance of separating LF and HF contributions to elicit the different roles of heart rate, arterial pressure and respiration on cardiovascular and cerebrovascular interactions. Moreover, the significant increase of redundancy with tilt suggests that these redundant effects are enhanced during postural stress, likely as a consequence of sympathetic activation and vagal withdrawal Malpas (2002); Porta et al. (2017). Therefore, the evidence that cardiovascular and cerebrovascular interactions occur through the coupling of rhythms in different frequency bands with different physiological meaning Porta and Faes (2015), make the proposed spectral PIRD eligible to probe high-order interactions in these networks Faes et al. (2021).
@@ Conclusions
Overall, by integrating temporal structure and frequency-specific dependencies, PIRD offers a more refined and comprehensive approach for studying multivariate interactions in neuroscience, biology, engineering, and beyond. Future works will explore further refinements in redundancy definitions and their applicability to real-world networked systems with nonlinear dependencies.
Acknowledgements.

### v3

@@ V.2 Data and statistical analysis
The four physiological time series described above were taken as a realization of the random processes F, M, H and R. These processes were investigated to decompose the rate of information shared dynamically between the target $Y=F$ and the set of source processes ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}=\{H,M,R\}$, with $H=X_{1},M=X_{2},R=X_{3}$, with the aim of investigating the mechanisms underlying the short-term dynamic regulation of the cerebral blood flow in resting and stress conditions. Each series was first detrended with an AR high-pass filter with zero phase (cutoff frequency 0.015 cycles/beat) Nollo et al. (2000). Then, a VAR model in the form of (31) was fitted to the four time series; model identification was performed via the ordinary least-squares approach Lütkepohl (2005), setting the model order $p$ according to the Akaike Information Criterion (AIC) for each subject (with maximum model order equal to 14) Akaike (1974b). After VAR identification, computation of time and frequency domain interaction measures of mutual, unique, redundant and synergistic information rates was performed from the estimated model parameters and spectra of the processes. Spectral analysis was performed assuming the series as uniformly sampled with the mean HP taken as the sampling period. Specifically, the spectral MIRs between the target and groups of sources were computed as in (28); then, according to coarse-graining PIRD, the unique, redundant and synergistic information rates calculated from the spectral redundancy rate in (23) were computed through integration within the low frequency (LF, $\left[0.04-0.15\right]$ Hz) and high frequency (HF, $\left[0.15-0.4\right]$ Hz) bands of the spectrum to analyze specific rhythms with physiological meaning Malpas (2002), as well as through integration over all frequencies, to get overall time-domain measures.
@@
Figure 7: Coarse-grained PIRD applied to the physiological network of mean cerebral blood flow velocity ($F$), mean arterial pressure ($M$), heart period ($H$) and respiration ($R$) assessed in patients prone to develop postural-related syncope. The mutual (left column: $I_{F;H}$ (black dots), $I_{F;M}$ (grey dots), $I_{F;R}$ (pink dots), $I_{F;H,M,R}$ (light blue dots)), unique (middle column: $\mathcal{U}_{F;H}$ (black dots), $\mathcal{U}_{F;M}$ (grey dots), $\mathcal{U}_{F;R}$ (pink dots)), redundant (right column: $\mathcal{R}_{F;H;M;R}$, magenta dots) and synergistic (right column: $\mathcal{S}_{F;H;M;R}$, blue dots) information rates shared between the target $F$ and the sources $\{H,M,R\}$ are computed in the RS (left boxplots), ET (middle boxplots) and LT (right boxplots) conditions along a) the whole frequency axis (TIME), b) the LF and c) the HF bands of the spectrum, taking the spectral redundancy rate function as in (23). The redundancy-synergy balance $\Delta$ is shown as boxplot distributions and individual values (purple dots) in the right column of panels a) (TIME), b) (LF) and c) (HF). The statistical differences among pairs of unique information rates (middle panels) were tested via Wilcoxon signed rank test for paired data ($p<0.05$) with Bonferroni-Holm correction for multiple comparisons: significant p-values are shown above the boxplot distributions with colors indicating the distribution to be compared with (i.e., black if compared with $\mathcal{U}_{F;H}$, grey if compared with $\mathcal{U}_{F;M}$). Wilcoxon signed rank test for paired data, $p<0.05$: $\ast$, RS vs. ET; §$\S$, RS vs. LT; ##\##, ET vs. LT.
@@ V.3 Results and discussion
Overall, these results confirm the redundant nature of cardiovascular and cerebrovascular interactions previously reported for similar triplets of physiological processes Faes et al. (2017); Porta et al. (2017); Faes et al. (2022), and document the relevance of separating LF and HF contributions to elicit the different roles of heart rate, arterial pressure and respiration on cardiovascular and cerebrovascular interactions. Moreover, the significant increase of redundancy with tilt suggests that these redundant effects are enhanced during postural stress, likely as a consequence of sympathetic activation and vagal withdrawal Malpas (2002); Porta et al. (2017). Therefore, the evidence that cardiovascular and cerebrovascular interactions occur through the coupling of rhythms in different frequency bands with different physiological meaning make the proposed spectral PIRD eligible to probe high-order interactions in these networks Faes et al. (2021).
@@ Conclusions
Overall, by integrating temporal structure and frequency-specific dependencies, PIRD offers a more refined and comprehensive approach for studying multivariate interactions in neuroscience, biology, engineering, and beyond. Future works will explore further refinements in redundancy definitions to be used in PIRD schemes involving purely discrete Williams and Beer (2010), continuous Ehrlich et al. (2024) and mixed Barà et al. (2024) random processes and different corresponding examples of dynamic interactions where to test them Timme et al. (2014); Barnett and Crutchfield (2015); Siggiridou et al. (2019), as well as investigate their applicability to real-world networked systems with nonlinear dependencies.
Appendix A
This appendix provides extended mathematical derivations to support and expand upon the new definition of spectral MMI redundancy rate discussed in Sect. "Frequency-domain PIRD". Derivations are given considering a dynamic network of $M+1$ random processes $\{Y,X_{1},\ldots,X_{M}\}$ where $Y$ is the target $\textbf{X}=\{X_{1},\ldots,X_{M}\}$ is the group of $M$ sources.
Appendix A.1
In this section, we demonstrate that the SMMI redundancy rate defined in (24) is less or equal than the MMI redundancy rate given in (17), with equality holding in the absence of temporal correlations.
Let us consider the generic $\alpha^{\mathrm{th}}$ atom of the lattice structure, $\alpha=\{\alpha_{1},\ldots,\alpha_{J}\}$, with cardinality $J$; the $j^{\mathrm{th}}$ element of the atom (i.e., $\alpha_{j}$) contains $K_{j}$ sources; e.g., the atom $\{\{1\}\{23\}\}$ has $J=2$ and $K_{1}=1,K_{2}=2$. We remark that the spectral redundancy rate function associated to the $\alpha^{\mathrm{th}}$ atom is computed at each frequency $\omega$ as in (23). Integration of the MIR spectral profiles $i_{Y;\textbf{X}_{\alpha_{j}}}(\omega)$, $j=1,\ldots,J$, in the range $\left[-\pi,\pi\right]$ returns the time-domain MIRs $I_{Y;\textbf{X}_{\alpha_{j}}}$ according to (18); the time-domain SMMI redundancy rate, $I^{\mathrm{SMMI}}_{Y;\textbf{X}_{\alpha}}$, is derived by (24). Moreover, taking the minimum of the set $\{I_{Y;\textbf{X}_{\alpha_{j}}}\}$, $j=1,\ldots,J$, yields the MMI redundancy rate $I^{\mathrm{MMI}}_{Y;\textbf{X}_{\alpha}}$ as in (17). In light of this, we demonstrate the following proposition.
Proposition A.1.1. $I^{\mathrm{SMMI}}_{Y;\textbf{X}_{\alpha}}\leq I^{\mathrm{MMI}}_{Y;\textbf{X}_{\alpha}}$.
Proof. Since the minimum spectral redundancy rate function is taken pointwise in the frequency domain, we have that
$i^{\mathrm{SMMI}}_{Y;\textbf{X}_{\alpha}}(\omega)\leq i_{Y;\textbf{X}_{\alpha_{j}}}(\omega)$
$\forall\omega\in\left[-\pi,\pi\right]$, $\forall j=1,\ldots,J$. If the functions $i_{Y;\textbf{X}_{\alpha_{j}}}(\omega)$ are non-negative and integrable in $\left[-\pi,\pi\right]$ (e.g., for Gaussian processes Nedungadi et al. (2011); Geweke (1982)), $i^{\mathrm{SMMI}}_{Y;\textbf{X}_{\alpha}}(\omega)$ is also non-negative and integrable in $\left[-\pi,\pi\right]$ Rudin (1976), resulting in
[…]
$\forall j=1,\ldots,J$, due to the property of monotonicity of the integral (i.e., the integral preserves order Rudin (1976)). This implies that
[…]
i.e.,
$I^{\mathrm{SMMI}}_{Y;\textbf{X}_{\alpha}}\leq\min_{j=1,\ldots,J}I_{Y,\textbf{X}_{\alpha_{j}}},$
ultimately demonstrating the stated proposition.
Equality holds in the absence of temporal correlations, i.e., when the random processes $\{Y,X_{1},\ldots,X_{M}\}$ reduce to i.i.d. random variables. If this is the case, the auto- and cross-correlation functions are Dirac delta functions, whose Fourier transforms return constant (cross-)spectral density matrices equal to the (co)variances of the processes, i.e.,
$P_{Y}(\omega)=\sigma^{2}_{Y};\quad\textbf{P}_{\textbf{X}_{\alpha_{j}}}(\omega)=\mathbf{\Sigma}_{\textbf{X}_{\alpha_{j}}};\quad\textbf{P}_{Y\textbf{X}_{\alpha_{j}}}(\omega)=\mathbf{\Sigma}_{Y;\textbf{X}_{\alpha_{j}}},$
where $\sigma^{2}_{Y}$ is the variance of the target, $\mathbf{\Sigma}_{\textbf{X}_{\alpha_{j}}}$ is the $K_{j}\times K_{j}$ auto-covariance matrix of $\textbf{X}_{\alpha_{j}}$ and $\mathbf{\Sigma}_{Y;\textbf{X}_{\alpha_{j}}}$ is the $1\times K_{j}$ cross-covariance of $Y$ and $\textbf{X}_{\alpha_{j}}$. Under the hypothesis of gaussianity, the MIR between the variables $Y$ and $\textbf{X}_{\alpha_{j}}$ reduces to $I_{Y;\textbf{X}_{\alpha_{j}}}=-\frac{1}{2}\log\big(1-\rho^{2}(Y;\textbf{X}_{\alpha_{j}})\big{missing})$, where
$\rho^{2}(Y;\textbf{X}_{\alpha_{j}})=\frac{\mathbf{\Sigma}_{Y;\textbf{X}_{\alpha_{j}}}\mathbf{\Sigma}^{-1}_{\textbf{X}_{\alpha_{j}}}\mathbf{\Sigma}^{\intercal}_{Y;\textbf{X}_{\alpha_{j}}}}{\sigma^{2}_{Y}}=1-\frac{|\mathbf{\Sigma}_{[Y\textbf{X}_{\alpha_{j}}]}|}{\sigma^{2}_{Y}|\mathbf{\Sigma}_{\textbf{X}_{\alpha_{j}}}|}$
is the squared correlation coefficient, with $\mathbf{\Sigma}_{[Y\textbf{X}_{\alpha_{j}}]}=\begin{pmatrix}\sigma^{2}_{Y}&\mathbf{\Sigma}_{Y;\textbf{X}_{\alpha_{j}}}\\ \mathbf{\Sigma}^{\intercal}_{Y;\textbf{X}_{\alpha_{j}}}&\mathbf{\Sigma}_{\textbf{X}_{\alpha_{j}}}\end{pmatrix}$ the full joint covariance matrix. This leads to
$I_{Y;\textbf{X}_{\alpha_{j}}}=\frac{1}{2}\log\biggr(\frac{\sigma^{2}_{Y}|\mathbf{\Sigma}_{\textbf{X}_{\alpha_{j}}}|}{|\mathbf{\Sigma}_{[Y\textbf{X}_{\alpha_{j}}]}|}\biggr{missing}),$
which yields
$I^{\mathrm{MMI}}_{Y;\textbf{X}_{\alpha}}=\min_{j=1,\ldots,J}I_{Y;\textbf{X}_{\alpha_{j}}}=\min_{j=1,\ldots,J}\frac{1}{2}\log\biggr(\frac{\sigma^{2}_{Y}|\mathbf{\Sigma}_{\textbf{X}_{\alpha_{j}}}|}{|\mathbf{\Sigma}_{[Y\textbf{X}_{\alpha_{j}}]}|}\biggr{missing}).$
In line with (29) and the statements above, the spectral MIR does not depend on the frequency $\omega$, displaying a flat profile $\forall\omega\in\left[-\pi,\pi\right]$, and can be written as
$i_{Y;\textbf{X}_{\alpha_{j}}}(\omega)=\frac{1}{2}\log\biggr(\frac{\sigma^{2}_{Y}|\mathbf{\Sigma}_{\textbf{X}_{\alpha_{j}}}|}{|\mathbf{\Sigma}_{[Y\textbf{X}_{\alpha_{j}}]}|}\biggr{missing});$
it follows that the spectral redundancy rate function $i^{\mathrm{SMMI}}_{Y;\textbf{X}_{\alpha}}$ does not vary with $\omega$, yielding
$I^{\mathrm{SMMI}}_{Y;\textbf{X}_{\alpha}}=\min_{j=1,\ldots,J}\frac{1}{2}\log\biggr(\frac{\sigma^{2}_{Y}|\mathbf{\Sigma}_{\textbf{X}_{\alpha_{j}}}|}{|\mathbf{\Sigma}_{[Y\textbf{X}_{\alpha_{j}}]}|}\biggr{missing}),$
ultimately demonstrating that $I^{\mathrm{MMI}}_{Y;\textbf{X}_{\alpha}}=I^{\mathrm{SMMI}}_{Y;\textbf{X}_{\alpha}}$.
Appendix A.2
In this section, we demonstrate that moving from the PIRD defined on the pointwise level (frequency-specific) to that defined on the process-level (time domain) is straightforward exploiting spectral integration and basic properties of the integral.
Proposition A.2.1. Solving the PIRD as
[…]
where, according to (21),
$i^{\delta}_{Y;\textbf{X}_{\alpha}}(\omega)=i^{\cap}_{Y;\textbf{X}_{\alpha}}(\omega)-\sum_{\beta\prec\alpha}i^{\delta}_{Y;\textbf{X}_{\beta}}(\omega),$
is equivalent to solve the PIRD as
[…]
where, according to (24),
[…]
Proof. The equivalence between statements (a) and (b) follows directly by applying the linearity of the integral Rudin (1976):
[…]
where $f(x)=i^{\cap}_{Y;\textbf{X}_{\alpha}}(\omega)$ and $g(x)=i^{\delta}_{Y;\textbf{X}_{\alpha}}(\omega)$. In this case, we have:
[…]
$=I^{\cap}_{Y;\textbf{X}_{\alpha}}-\sum_{\beta\prec\alpha}I^{\delta}_{Y;\textbf{X}_{\beta}},$
which confirms that the frequency-domain decomposition (a) and the time-domain decomposition (b) are equivalent.
Appendix A.3
In this section, we demonstrate that the SMMI spectral redundancy rate defined by (23, 29) satisfies the properties of weak symmetry (Proposition A.3.1.), self-redundancy (Proposition A.3.2.), monotonicity (Proposition A.3.3.) and non-negativity (Proposition A.3.4.) proposed by Williams and Beer Williams and Beer (2010).
Proposition A.3.1. (weak symmetry) The SMMI spectral redundancy rate is invariant with the ordering of the sources, i.e., $i^{\mathrm{SMMI}}_{Y;\textbf{X}_{\{\alpha_{1},\ldots,\alpha_{J}\}}}(\omega)$ is symmetric w.r.t. the $\alpha_{j},j=1,\ldots,J$.
Proof. The proposition follows immediately from the property of commutativity of the minimum operator applied to (23).
Proposition A.3.2. (self-redundancy) The SMMI spectral redundancy rate applied to a single collection of sources reduces to the spectral MIR between the sources and the target, i.e. $i^{\mathrm{SMMI}}_{Y;\textbf{X}_{\alpha}}(\omega)=i_{Y;\textbf{X}_{\alpha}}(\omega)$ when $J=|\alpha|=1$.
Proof. The proposition follows immediately from (23) applied with $J=|\alpha|=1$.
Proposition A.3.3. (monotonicity and subset equality) The SMMI spectral redundancy rate does not increase when adding sources, i.e., $i^{\mathrm{SMMI}}_{Y;\textbf{X}_{\{\alpha_{1},\ldots,\alpha_{J-1},\alpha_{J}\}}}(\omega)\leq i^{\mathrm{SMMI}}_{Y;\textbf{X}_{\{\alpha_{1},\ldots,\alpha_{J-1}\}}}(\omega)$, with equality if $\alpha_{J-1}\subseteq\alpha_{J}$.
Proof. Let us denote:
  • $\alpha^{\prime}=\{\alpha_{1},\ldots,\alpha_{J-1}\}$,
  • $\alpha^{\prime\prime}=\alpha\cup\{\alpha_{J}\}=\{\alpha_{1},\ldots,\alpha_{J-1},\alpha_{J}\}$.
The spectral redundancy rate function over the expanded collection is given by:
$i^{\cap}_{Y;\textbf{X}_{\alpha^{\prime\prime}}}(\omega)=\min\{i_{Y;\textbf{X}_{\alpha_{1}}}(\omega),\ldots,\\ i_{Y;\textbf{X}_{\alpha_{J-1}}}(\omega),i_{Y;\textbf{X}_{\alpha_{J}}}(\omega)\}.$
Similarly, the spectral redundancy rate function over the original collection is:
$i^{\cap}_{Y;\textbf{X}_{\alpha^{\prime}}}(\omega)=\min\{i_{Y;\textbf{X}_{\alpha_{1}}}(\omega),\ldots,i_{Y;\textbf{X}_{\alpha_{J-1}}}(\omega)\}.$
Since taking the minimum over a larger set of values can only decrease or preserve the original minimum value, it follows that
$i^{\cap}_{Y;\textbf{X}_{\alpha^{\prime\prime}}}(\omega)\leq i^{\cap}_{Y;\textbf{X}_{\alpha^{\prime}}}(\omega),$
thus proving monotonicity.
Subset equality, $i^{\cap}_{Y;\textbf{X}_{\alpha^{\prime\prime}}}(\omega)=i^{\cap}_{Y;\textbf{X}_{\alpha^{\prime}}}(\omega)$ if $\alpha_{J-1}\subseteq\alpha_{J}$, holds thanks to the use of the minimum operator if
$i_{Y;\textbf{X}_{\alpha_{J}}}(\omega)\geq i_{Y;\textbf{X}_{\alpha_{J-1}}}(\omega);$
the latter follows from the application of the chain rule to the spectral MIR $i_{Y;\textbf{X}_{\alpha_{J}}}(\omega)$. Specifically, denoting $\alpha_{J-1}=\{\alpha_{J-1}^{1},\ldots,\alpha_{J-1}^{K_{J-1}}\}$ and $\alpha_{J}=\{\alpha_{J-1},\alpha_{J}^{K_{{J-1}}+1},\ldots,\alpha_{J}^{K_{J}}\}$, the chain rule reads
$i_{Y;\textbf{X}_{\alpha_{J}}}(\omega)=i_{Y;\textbf{X}_{\alpha_{J-1}}}+i_{Y;\textbf{X}_{\alpha_{J}^{{K_{{J-1}}+1}}},\ldots,\textbf{X}_{\alpha_{J}^{{K_{J}}}}|\textbf{X}_{\alpha_{J-1}}}(\omega),$
which proves that $i_{Y;\textbf{X}_{\alpha_{J}}}(\omega)\geq i_{Y;\textbf{X}_{\alpha_{J-1}}}(\omega)$ given the non-negativity of $i_{Y;\textbf{X}_{\alpha_{J}^{{K_{{J-1}}+1}}},\ldots,\textbf{X}_{\alpha_{J}^{{K_{J}}}}|\textbf{X}_{\alpha_{J-1}}}(\omega)$.
Proposition A.3.4. (non-negativity) The SMMI spectral redundancy rate is non negative, i.e. $i^{\mathrm{SMMI}}_{Y;\textbf{X}_{\alpha}}(\omega)\geq 0$ $\forall\omega\in[-\pi,\pi]$.
Proof. The proof is given showing that each spectral MIR defined in (29) at varying $j=1,\ldots,J$ is non-negative. Applying the generalized Hadamard’s inequality Friedland (1975) to the spectral density matrix (30) yields
$|\textbf{P}_{[Y\textbf{X}_{\alpha_{j}}]}(\omega)|\leq P_{Y}(\omega)|\textbf{P}_{\textbf{X}_{\alpha_{j}}}(\omega)|,$
which corresponds to
$i_{Y;{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}}}(\omega)=\frac{1}{2}\log\frac{|\textbf{P}_{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\textbf{X}}_{\alpha_{j}}}(\omega)|P_{Y}(\omega)}{|\textbf{P}_{[Y\textbf{X}_{\alpha_{j}}]}(\omega)|}\geq 0,$
thus completing the proof.
Acknowledgments